The function C(x) = 119x +750 represents the cost of producing x three-pronged stainless steel widgets. R(x) =169x represents the revenue from selling x of them. Find the following
a. Number that must be sold to break even
b. Profit from selling 500
c. Profit or loss from selling 200 at the regular price and 300 at half price.

Answer by ankor@dixie-net.com(22740)

The function C(x) = 119x +750 represents the cost of producing x three-pronged stainless steel widgets. R(x) =169x represents the revenue from selling x of them. Find the following:
:
a. Number that must be sold to break even
This occurs when R(x) + C(x), therefore
169x = 119x + 750
169x - 119x = 750
50x = 750
x = 750/50
x = 15 widgets to break even
:
b. Profit from selling 500
Profit is revenue - costs:
P(x) = 169x - (119x+750)
x = 500
P(x) = 169(500) - (119(500) + 750)
P(x) = 84500 - 59500 - 750
P(x) = $24,250 is the profit from selling 500
:
c. Profit or loss from selling 200 at the regular price and 300 at half price.
Regular price is $169 per unit, half price: $84.50
find the revenue from 200 units at regular price: 200*169 = $33,800
find revenue from 300 units sold at half price: 300*84.50 = $25,350
------------------------------------------------------------------------
find the total revenue selling 500 units at these prices: $59,150
:
Find the total cost of making these 500 units: 500(119)+750=$60,250
59,150 - 60250 = -$1100 loss
:
This is the method, check my math.